OpenStudy (anonymous):

Choose the best description for the real number 4 point 6 repeating. A. Irrational, because it is a terminating decimal B. Irrational, because it is a repeating decimal C. Rational, because it is a terminating decimal D. Rational, because it is a repeating decimal

OpenStudy (anonymous):

thank you @steve816

OpenStudy (anonymous):

let x=4.666.... (1) 10x=46.666... (2) 10x-x=42 9x=42 \[x=\frac{ 42 }{ 9 }=\frac{ 14 }{ 3 }=?\]

OpenStudy (anonymous):

what can you say now?